OpenStudy (anonymous):

Consider the following sequence: \[\sqrt{2},\sqrt{2\sqrt{2}}\,\sqrt{2\sqrt{2\sqrt{2}}}\] Prove that it converges, and find it's limit. EDIT: Sorry, wrong question

OpenStudy (kc_kennylau):

Find \(\Large\displaystyle\lim_{n\rightarrow\infty}\frac{a_{n+1}}{a_n}\)

OpenStudy (anonymous):

well, did you find the general equation for a(n) ?

OpenStudy (anonymous):

@sourwing my guess is: a(n)=sqrt(n sqrt(a(n-1))) with a(1)=sqrt(n)

OpenStudy (kc_kennylau):

Bear in mind that \(\sqrt n=n^{0.5}\)

OpenStudy (anonymous):

well, that's the recursive formula. I was talking about explicit formula

OpenStudy (anonymous):

how about let's just start it. a(1) = 2^(1/2) a(2) = 2^(1/2) * 2^(1/4) a(3) = 2^(1/2) * 2^(1/4) * 2^(1/8) see the pattern?

OpenStudy (anonymous):

yes I see the pattern

OpenStudy (anonymous):

so what do you think a(n) is?

OpenStudy (anonymous):

I don't know how to state it as a explicit formula, as I don't have much experience with sequences. However, I can clearly see the pattern and state the recursive formula as I did above.

OpenStudy (kc_kennylau):

@sourwing don't let him mess with \(\Large\displaystyle\Pi\)

OpenStudy (anonymous):

that's fine. These type of problems aren't easy to begin with anyways. notice that: a(3) = 2^(1/2 + 1/4 + 1/8) using exponent rule then a(n) = 2^(1/2 + 1/4 + 1/8 + ... + 1/2^n) yes? @kc_kennylau i'm not :D

OpenStudy (anonymous):

@kc_kennylau as a matter of fact, i'm a noob when it comes to pi product XD

OpenStudy (anonymous):

yes, i see it now

OpenStudy (anonymous):

good, now can you write it in terms of summation notation?

OpenStudy (anonymous):

we haven't talked about summation in class, but would be my guess: \[\sum_{i=m}^{n} 2^{1/2^{m}}+2^{1/2^{m+1}}....2^{1/2^{n}}\]

OpenStudy (anonymous):

I meant to say: \[\sum_{i=m}^{n}2^{1/2^i}\]

OpenStudy (anonymous):

you can't do that because 2^(a + b) isn't 2^a + 2^b. the summation is on the exponent only. so \[a(n) = 2^{\sum_{i=1}^{n} (1/2)^i}\] yes?

OpenStudy (anonymous):

that's a letter i in case you can't read it

OpenStudy (anonymous):

ya that makes sense, I see what I did wrong

OpenStudy (anonymous):

Now, technically this is only a propose for the formula of a(n). Even though it's obviously true, we still need to prove that is true for all natural n.

OpenStudy (anonymous):

any idea which method of proving we are going to use?

OpenStudy (anonymous):

induction

OpenStudy (anonymous):

bingo! suppose it's true for a(k) for some k, then:  yes?